# Simple homotopy types of even dimensional manifolds

### v2

@@ 7. The proof of Theorems B, C, LABEL:, and D
(a), (b), (d), (e), (g), and (h) follow immediately from Lemma 7.2 and Theorems 5.14, 5.15, LABEL:, and 5.17.
(c) For every $m$, it follows from Lemma 7.2 (c) that $|\mathcal{M}^{h}_{s}(M^{n}_{m})|>\frac{|\mathcal{J}_{n}(C_{\infty}\times C_{m})|}{2m^{2}}$ for every $n$. By Theorem 5.14 (iii) $\frac{|\mathcal{J}_{n}(C_{\infty}\times C_{m})|}{2m^{2}}\to\infty$ as $m\to\infty$. Item (f), can be proved similarly using Theorem 5.15 (iii).
(i) For every $m$, it follows from Lemma 7.2 (i) that $\inf_{n}|\mathcal{M}^{h}_{s,\operatorname{hCob}}(M^{n}_{m})|>\frac{|\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))|}{2m^{2}}$. By Theorem 5.17 (iii), we see that $\frac{1}{2m^{2}}\sup_{l\leq m}|\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{l}))|\to\infty$ as $m\to\infty$. This implies that $\sup_{l\leq m}\frac{|\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{l}))|}{2l^{2}}\to\infty$ as $m\to\infty$. So the map $m\mapsto\frac{|\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))|}{2m^{2}}$ is unbounded. Therefore $\limsup_{m\to\infty}\frac{|\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))|}{2m^{2}}=\infty$. ∎
This completes the proof of the remaining parts of Theorems B, C, LABEL:, and D, subject to the proofs of Theorems 11.1, 11.2, and 11.3 which will be postponed until Section 11. We can now deduce:
Theorem 7.4.
Proof.
By Theorem 7.3 (b), if $m$ is not square-free (e.g. if $m=4$), then $\mathcal{M}^{h}_{s}(M^{n}_{m})$ is infinite. Note that $M^{n}_{m}=S^{1}\times L$ is orientable, so the same is true for every $n$-manifold homotopy equivalent to it. Therefore we can get a suitable infinite collection by choosing a representative from each element of $\mathcal{M}^{h}_{s}(M^{n}_{m})$. ∎
This completes the proof of Theorem A. Note that this does not depend on the results postponed to Section 11. In particular, if $m$ is not square-free, then $|NK_{1}(\mathbb{Z}C_{m})|=\infty$ (Theorem 1.5) which implies $|\mathcal{J}_{n}(C_{\infty}\times C_{m})|=\infty$ (Lemma 5.12) and so $|\mathcal{M}^{h}_{s}(M^{n}_{m})|=\infty$ (Lemma 7.2 (b)).
@@
The aim of this part will be to prove Theorems 11.1, 11.2, LABEL:, and 11.3, which are key ingredients in the proofs of Theorems 5.14, 5.15, and 5.17 respectively. In Sections 8 and 9, we will recall the necessary background on class groups, Tate cohomology, and $\mathbb{Z}C_{2}$-modules. The main technical heart of this part will be Section 10 where we will investigate the involution on $\widetilde{K}_{0}(\mathbb{Z}C_{m})$ and prove general results which allow it to be computed. In Section 11, we make use of the results in Section 10 to prove Theorems 11.1, 11.2, LABEL:, and 11.3. Throughout, we assume that all modules are left modules.
@@ 8. Locally free class groups
We will now recall the theory of locally free class groups for orders in semisimple $\mathbb{Q}$-algebras. Good references for this material are [Sw80, Section 1-3] and [CR87, Section 49A & 50E].
@@ 8.1. Definitions and properties
Recall that for a ring $A$, a nonzero $A$-module is simple if it contains no simple $A$-submodules other than itself and 00, and is semisimple if isomorphic as an $A$-module to a direct sum of its simple $A$-submodules. We say that a ring $A$ is simple (resp. semisimple) if $A$, viewed as an $A$-module, is simple (resp. semisimple). For a field $K$, a $K$-algebra is a ring $A$ for which $K$ is a subring of the centre $Z(A)$. A $K$-algebra is finite-dimensional if it is finite-dimensional as a $K$-vector space.
@@
A $\Lambda$-module $M$ is locally free if $M_{p}=\mathbb{Z}_{p}\otimes_{\mathbb{Z}}M$ is a free $\Lambda_{p}$-module for all primes $p$.
The following is [Sw80, Lemma 2.1]. Note that the converse need not hold, i.e. there exist orders $\Lambda$ and projective $\Lambda$-modules that are not locally free [Sw80, p. 156].
@@
It follows that $C(\Lambda)\leq\widetilde{K}_{0}(\Lambda)$ is a subgroup, where $\widetilde{K}_{0}$ is the 00th reduced algebraic $K$-group as defined in Section 5.1. It is a consequence of the Jordan-Zassenhaus theorem that $C(\Lambda)$ is finite [CR87, Remark 49.11 (ii)].
We will now specialise to the case where $\Lambda=\mathbb{Z}G$ for $G$ a finite group. In contrast to the situation for general orders, we have the following [Sw80, p. 156]. By Proposition 8.2, this implies that a $\mathbb{Z}G$-module is projective if and only if it is locally free.
@@
    If $M$ is a projective $\mathbb{Z}G$-module, then $M$ is locally free.
@@
Finally we note the following which relates locally free class groups to ideal class groups (see [Re75, Section 35]).
@@ 8.2. Kernel groups
Let $A$ be a finite-dimensional semisimple $\mathbb{Q}$-algebra $A$. An order in $A$ is said to be maximal if it is not properly contained in another other order in $A$. Since every finite field extension of $\mathbb{Q}$ is separable, $A$ is a separable algebra and so every order in $A$ is contained in a maximal order [Sw70, Proposition 5.1].
Let $\Lambda$ be an order in $A$ and let $\Gamma$ be a maximal order in $A$ containing $\Lambda$. The inclusion map $i\colon\Lambda\hookrightarrow\Gamma$ induces a map $i_{*}\colon C(\Lambda)\to C(\Gamma)$ given by extension of scalars $[M]\mapsto[\Gamma\otimes_{\Lambda}M]$ which is necessarily surjective by [CR87, Theorem 49.25].
@@
Define the kernel group $D(\Lambda)$ to be the kernel of the map $i_{*}\colon C(\Lambda)\to C(\Gamma)$. This is often also referred to as the defect group.
The group $C(\Gamma)$ does not depend on the choice of maximal order in $A$; in fact, if $\Gamma_{1}$, $\Gamma_{2}$ are maximal orders in $A$, then $C(\Gamma_{1})\cong C(\Gamma_{2})$ [CR87, Theorem 49.32]. Furthermore, the kernel group $D(\Lambda)$ does not depend on the choice of maximal order. This can be seen from the fact that it can be defined without reference to a maximal order: if $M$ is a locally free $\Lambda$-module, then $[M]\in D(\Lambda)$ if and only if there exists a finitely generated $\Lambda$-module $X$ such that $M\oplus X\cong\Lambda^{n}\oplus X$ for some $n$ [CR87, Proposition 49.34].
@@ 8.3. The idèlic approach to locally free class groups
As a subgroup of $A_{p}^{\times}$, this is independent of the choice of order $\Lambda$ [CR87, p. 218]. Every class in $C(\Lambda)$ is represented by a locally free $\Lambda$-module $M\subseteq A$ [CR87, p. 218]. For each $p$, there exists $\alpha_{p}\in A_{p}$ such that $M_{p}=\Lambda_{p}\alpha_{p}\subseteq A_{p}$. For all but finitely many $p$, $M_{p}\cong\Lambda_{p}$ and so $\alpha_{p}\in\Lambda_{p}^{\times}$. In particular, $\alpha=(\alpha_{p})\in J(A)$. Conversely, given an idèle $\alpha\in J(A)$, we have that $\Lambda\alpha=A\cap\bigcap_{p}\Lambda_{p}\alpha_{p}\subseteq A$ is a locally free $\Lambda$-ideal. Let $\alpha,\beta\in J(A)$. Then $\Lambda\alpha\cong\Lambda\beta$ as $\Lambda$-modules if and only if $\beta\in U(\Lambda)\cdot\alpha\cdot A^{\times}$ where $A^{\times}\subseteq J(A)$ by sending $a\in A^{\times}$ to $\alpha_{p}=1\otimes_{\mathbb{Q}}a$ for all $p$, and $U(\Lambda)=\{(\alpha_{p})\in\prod_{p}A_{p}^{\times}\mid\alpha_{p}\in\Lambda_{p}^{\times}\text{ for all $p$}\}\subseteq J(A)$ [CR87, 49.6]. Furthermore, we have that $\Lambda\alpha\oplus\Lambda\beta\cong\Lambda\oplus\Lambda\alpha\beta$ [CR87, 49.8]. This leads to the following.
@@
$[\Lambda\,\cdot\,]\colon J(A)\twoheadrightarrow C(\Lambda),\quad\alpha\mapsto[\Lambda\alpha].$
@@
Whilst we will not make use of it in this article, we note that this leads to the formula
@@ 8.4. Involutions on locally free class groups
Let $\Lambda$ be an order in a finite-dimensional semisimple $\mathbb{Q}$-algebra $A$. Suppose further that $A$ is a ring equipped with an involution ¯$\overline{\cdot}\colon A\to A$ which restricts to $\Lambda$, i.e. the map $\overline{\cdot}$ is an involution on $A$ as an abelian group which satisfies $\overline{xy}=\overline{y}\cdot\overline{x}$ for all $x,y\in A$ and $\overline{x}\in\Lambda$ for all $x\in\Lambda$. For example, if $G$ is a finite group and $w\colon G\to\{\pm 1\}$ is an orientation character, then $A=\mathbb{Q}G$ has an involution given by $\sum_{i=1}^{k}n_{i}g_{i}\mapsto\sum_{i=1}^{k}w(g_{i})n_{i}g_{i}^{-1}$ for $n_{i}\in\mathbb{Z}$ and $g_{i}\in G$ which restricts to $\Lambda=\mathbb{Z}G$ (see Section 2.2). Given a (left) $\Lambda$-module $M$, we now have the notion of a dual (left) $\Lambda$-module $M^{*}$ (see Definition 2.23).
We now note the following properties of the dual of locally free modules. This follows from the fact that, by Proposition 8.2, locally free modules are projective and hence reflexive.
@@
Let $M$ be a locally free $\Lambda$-module. Then:
  1. (i)
    $M^{*}$ is locally free;
  2. (ii)
    The evaluation map $\text{ev}:M\to M^{**}$, $m\mapsto(f\mapsto f(m))$ is an isomorphism of $\Lambda$-modules.
@@
There is an involution of abelian groups:
@@
That is, $\ast$ is a group homomorphism such that $\ast^{2}=\operatorname{Id}_{C(\Lambda)}$.
In the case $\Lambda=\mathbb{Z}G$, we have that $\widetilde{K}_{0}(\mathbb{Z}G)\cong C(\mathbb{Z}G)$ and the involution above coincides with the standard involution on $\widetilde{K}_{0}$ as defined in Section 5.1.
We will now explore some properties of this involution. In what follows we refer to [CR87, pp. 275-6]. This deals only with the case $\Lambda=\mathbb{Z}G$, though the arguments there apply to the more general setting described above.
It is an immediate consequence of the alternative description of the kernel group given in Section 8.2 that, if $[M]\in D(\Lambda)$, then $[M^{*}]\in D(\Lambda)$ [CR87, p. 275]. In particular, the involution $\ast$ restricts to $D(\Lambda)$. This implies that, if $i\colon\Lambda\hookrightarrow\Gamma$ for $\Gamma$ a maximal order in $A$, then $\ast$ induces an involution on $C(\Gamma)$ via the map $i_{*}\colon C(\Lambda)\twoheadrightarrow C(\Gamma)$, and this coincides with the involution on $C(\Gamma)$ coming from the fact that $\Gamma$ is an order in $A$.
Recall that an involution on an abelian group is the same structure as a $\mathbb{Z}C_{2}$-module, where the $C_{2}$-action is given by the involution. In particular, we have shown the following.
@@
There is a short exact sequence of $\mathbb{Z}C_{2}$-modules
$0\to D(\Lambda)\to C(\Lambda)\to C(\Gamma)\to 0$
where $D(\Lambda)$, $C(\Lambda)$, and $C(\Gamma)$ are $\mathbb{Z}C_{2}$-modules under the involutions described above.
We will conclude this section by noting that the idèlic approach to class groups gives a different way to define an involution on $C(\Lambda)$. The involution ¯$\overline{\cdot}\colon A\to A$ induces involutions on $A_{p}$ for each $p$ and so on $J(A)$. It can be shown that the involution fixes the subgroups $A^{\times}$, $U(\Lambda)$ and $J_{0}(A)$ and so induces an involution on $C(\Lambda)$ given by $[\Lambda\alpha]\mapsto[\Lambda\overline{\alpha}]$ (see [CR87, p. 274]).
@@
The involution on $C(\Lambda)$ induced by the involution on $J(A)$ is the standard involution $\ast\colon C(\Lambda)\to C(\Lambda)$, $[M]\mapsto-[M^{*}]$.
This gives an alternate way to understand the standard involution on $\widetilde{K}_{0}(\mathbb{Z}G)$. We make further use of this description the following section.

### v3

@@ 7. The proof of Theorems A, C, D, LABEL:, and E
(a), (b), (d), (e), (g), and (h) follow immediately from Lemma 7.3 and Theorems 5.14 and 5.15.
(c) For every $m$, it follows from Lemma 7.3 (c) that $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M^{n}_{m})|>\frac{|\mathcal{J}_{n}(C_{\infty}\times C_{m})|}{2m^{2}}$ for every $n$. By Theorem 5.14 (iv) $\frac{|\mathcal{J}_{n}(C_{\infty}\times C_{m})|}{2m^{2}}\to\infty$ as $m\to\infty$. Item (f) can be proved similarly.
(i) For every $m$, it follows from Lemma 7.3 (i) that $\inf_{n}|\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M^{n}_{m})|>\frac{|\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))|}{2m^{2}}$. By Theorem 5.15 (iii), we see that $\frac{1}{2m^{2}}\sup_{l\leq m}|\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{l}))|\to\infty$ as $m\to\infty$. This implies that $\sup_{l\leq m}\frac{|\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{l}))|}{2l^{2}}\to\infty$ as $m\to\infty$. So the map $m\mapsto\frac{|\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))|}{2m^{2}}$ is unbounded. Therefore $\limsup_{m\to\infty}\frac{|\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))|}{2m^{2}}=\infty$. ∎
@@
The aim of this part is to prove Theorems 11.1, LABEL:, and 11.2, which are key ingredients in the proofs of Theorems 5.14 and 5.15 respectively. In Sections 8 and 9, we recall the necessary background on class groups, Tate cohomology, and $\mathbb{Z}C_{2}$-modules. The main technical heart of this part is Section 10 where we investigate the involution on $\widetilde{K}_{0}(\mathbb{Z}C_{m})$ and prove general results which allow it to be computed. In Section 11, we make use of the results in Section 10 to prove Theorems 11.1 and 11.2. Throughout, we assume that all modules are left modules.
@@ 8. Locally free class groups
We now recall the theory of locally free class groups for orders in semisimple $\mathbb{Q}$-algebras. Good references for this material are [Sw80, Section 1-3] and [CR87, Section 49A & 50E].
@@ 8.1. Definitions and properties
For a ring $A$, a nonzero $A$-module is simple if it contains no $A$-submodules other than itself and 0, and is semisimple if it is isomorphic as an $A$-module to a direct sum of its simple $A$-submodules. We say that a ring $A$ is semisimple if $A$, viewed as an $A$-module, is semisimple.
@@
A $\Lambda$-module $M$ is locally free if it is finitely generated and $M_{p}=\mathbb{Z}_{p}\otimes_{\mathbb{Z}}M$ is a free $\Lambda_{p}$-module for all primes $p$.
The following is [Sw80, Lemma 2.1]. The converse need not hold, i.e. there exist orders $\Lambda$ and finitely generated projective $\Lambda$-modules that are not locally free [Sw80, p. 156].
@@
It follows that $C(\Lambda)\leq\widetilde{K}_{0}(\Lambda)$ is a subgroup, where $\widetilde{K}_{0}$ is the 0th reduced algebraic $K$-group as defined in Section 5.1. It is a consequence of the Jordan-Zassenhaus theorem that $C(\Lambda)$ is finite [CR87, Remark 49.11 (ii)].
We now specialise to the case where $\Lambda=\mathbb{Z}G$ for $G$ a finite group. In contrast to the situation for general orders, we have the following proposition [Sw80, p. 156]. By Proposition 8.2, this implies that a finitely generated $\mathbb{Z}G$-module is projective if and only if it is locally free.
@@
    If $M$ is a finitely generated projective $\mathbb{Z}G$-module, then $M$ is locally free.
@@
Finally we note the following which relates locally free class groups to ideal class groups (see [Re75, Section 35]). For a field $K$, we let $\mathcal{O}_{K}$ denote the ring of integers.
@@ 8.2. Kernel groups
Let $A$ be a finite-dimensional semisimple $\mathbb{Q}$-algebra. An order in $A$ is said to be maximal if it is not properly contained in another order in $A$. Since every finite field extension of $\mathbb{Q}$ is separable, $A$ is a separable algebra and so every order in $A$ is contained in a maximal order [Sw70, Proposition 5.1].
Let $\Lambda$ be an order in $A$ and let $\Gamma$ be a maximal order in $A$ containing $\Lambda$. The inclusion map $i\colon\Lambda\hookrightarrow\Gamma$ induces a map $i_{*}\colon C(\Lambda)\to C(\Gamma)$ given by extension of scalars $M]\mapsto[\Gamma\otimes_{\Lambda}M]$ which is necessarily surjective by [CR87, Theorem 49.25].
@@
Define the kernel group $D(\Lambda)$ to be the kernel of the map $i_{*}\colon C(\Lambda)\to C(\Gamma)$. (This is often also referred to as the defect group.)
The group $C(\Gamma)$ does not depend up to isomorphism on the choice of maximal order in $A$, i.e. if $\Gamma_{1}$, $\Gamma_{2}$ are maximal orders in $A$ containing $\Lambda$, then there is an isomorphism $C(\Gamma_{1})\cong C(\Gamma_{2})$ [CR87, Theorem 49.32]. Furthermore, the kernel group $D(\Lambda)$ does not depend on the choice of maximal order. This can be seen from the fact that it can be defined without reference to a maximal order: if $M$ is a locally free $\Lambda$-module, then $M]\in D(\Lambda)$ if and only if there exists a finitely generated $\Lambda$-module $X$ such that $M\oplus X\cong\Lambda^{n}\oplus X$ for some $n$ [CR87, Proposition 49.34].
@@ 8.3. The idèlic approach to locally free class groups
As a subgroup of $\prod_{p}A_{p}^{\times}$, this is independent of the choice of order $\Lambda$ [CR87, p. 218]. Every class in $C(\Lambda)$ is represented by a locally free $\Lambda$-module $M\subseteq A$ [CR87, p. 218]. For each $p$, there exists $\alpha_{p}\in A_{p}$ such that $M_{p}=\Lambda_{p}\alpha_{p}\subseteq A_{p}$. For all but finitely many $p$, $M_{p}\cong\Lambda_{p}$ and so $\alpha_{p}\in\Lambda_{p}^{\times}$. In particular, $\alpha=(\alpha_{p})\in J(A)$. Conversely, given an idèle $\alpha\in J(A)$, we have that $\Lambda\alpha=A\cap\bigcap_{p}\Lambda_{p}\alpha_{p}\subseteq A$ is a locally free $\Lambda$-ideal. Let $\alpha,\beta\in J(A)$. Then $\Lambda\alpha\cong\Lambda\beta$ as $\Lambda$-modules if and only if $\beta\in U(\Lambda)\cdot\alpha\cdot A^{\times}$ where $A^{\times}\subseteq J(A)$ by sending $a\in A^{\times}$ to $\alpha_{p}=1\otimes_{\mathbb{Q}}a$ for all $p$, and $U(\Lambda)=\{(\alpha_{p})\in\prod_{p}A_{p}^{\times}\mid\alpha_{p}\in\Lambda_{p}^{\times}\text{ for all $p$}\}\subseteq J(A)$ [CR87, 49.6]. Furthermore, we have that $\Lambda\alpha\oplus\Lambda\beta\cong\Lambda\oplus\Lambda\alpha\beta$ [CR87, 49.8]. This leads to the following.
@@
$\Lambda\,\cdot\,]\colon J(A)\twoheadrightarrow C(\Lambda),\quad\alpha\mapsto[\Lambda\alpha].$
@@
This leads to the formula
@@ 8.4. Involutions on locally free class groups
Let $\Lambda$ be an order in a finite-dimensional semisimple $\mathbb{Q}$-algebra $A$. Suppose further that $A$ is a ring equipped with an involution $\overline{\cdot}\colon A\to A$ which preserves $\Lambda$, i.e. the map $\overline{\cdot}$ is an involution on $A$ as an abelian group which satisfies $\overline{xy}=\overline{y}\cdot\overline{x}$ for all $x,y\in A$ and $\overline{x}\in\Lambda$ for all $x\in\Lambda$. For example, if $G$ is a finite group and $w\colon G\to\{\pm 1\}$ is a homomorphism, then $A=\mathbb{Q}G$ has an involution given by $\sum_{i=1}^{k}n_{i}g_{i}\mapsto\sum_{i=1}^{k}w(g_{i})n_{i}g_{i}^{-1}$ for $n_{i}\in\mathbb{Z}$ and $g_{i}\in G$ which preserves $\Lambda=\mathbb{Z}G$ (see Section 2.2). Given a (left) $\Lambda$-module $M$, we now have the notion of a dual (left) $\Lambda$-module $M^{*}$ (see Definition 2.20).
We now note the following properties of the dual of locally free modules. An $R$-module $M$ is reflexive if the evaluation map $\text{ev}\colon M\to M^{**}$, $m\mapsto(f\mapsto f(m))$ is bijective.
@@
If $M$ is a locally free $\Lambda$-module, then $M$ is reflexive and $M^{*}$ is locally free.
Proof.
By Proposition 8.2, locally free modules are projective. It is well known that finitely generated projective modules are reflexive, hence $M$ is reflexive. The fact that $M^{*}$ is locally free follows from that fact that $\mathbb{Z}_{p}\otimes_{\mathbb{Z}}M^{*}\cong(\mathbb{Z}_{p}\otimes_{\mathbb{Z}}M)^{*}\cong\Lambda_{p}^{*}\cong\Lambda_{p}$ for all primes $p$. ∎
@@
Define the standard involution on $C(\Lambda)$ to be the map:
@@
It can be shown that this is an involution of abelian groups in the sense that $\ast$ is a group homomorphism such that $\ast^{2}=\operatorname{Id}_{C(\Lambda)}$. In the case $\Lambda=\mathbb{Z}G$, we have that $\widetilde{K}_{0}(\mathbb{Z}G)\cong C(\mathbb{Z}G)$ and the involution above coincides with the standard involution on $\widetilde{K}_{0}$ as defined in Section 5.1.
We now explore some properties of this involution. In what follows we refer to [CR87, pp. 275-6]. This deals only with the case $\Lambda=\mathbb{Z}G$, though the arguments there apply to the more general setting described above. Observe that an involution on an abelian group is the same structure as a $\mathbb{Z}C_{2}$-module, where the $C_{2}$-action is given by the involution.
@@
Let $\Gamma$ be a maximal order in $A$ containing $\Lambda$ and let $i\colon\Lambda\hookrightarrow\Gamma$ denote the inclusion map. Then there is a short exact sequence of $\mathbb{Z}C_{2}$-modules
$0\to D(\Lambda)\hookrightarrow C(\Lambda)\xrightarrow[]{i_{*}}C(\Gamma)\to 0$
where $D(\Lambda)$, $C(\Lambda)$, and $C(\Gamma)$ are $\mathbb{Z}C_{2}$-modules under the standard involutions.
Proof.
It is an immediate consequence of the alternative description of the kernel group given in Section 8.2 that, if $M]\in D(\Lambda)$, then $M^{*}]\in D(\Lambda) [$CR87, p. 275]. In particular, the involution $\ast$ restricts to $D(\Lambda)$, and the map $D(\Lambda)\hookrightarrow C(\Lambda)$ is a $\mathbb{Z}C_{2}$-module homomorphism. Thus $\ast$ induces an involution on $C(\Gamma)$ via the map $i_{*}$, and this coincides with the involution on $C(\Gamma)$ coming from the fact that $\Gamma$ is an order in $A$. ∎
We conclude this section by noting that the idèlic approach to class groups gives a different way to define an involution on $C(\Lambda)$. The involution $\overline{\cdot}\colon A\to A$ induces involutions on $A_{p}$ for each $p$ and so on $J(A)$. It can be shown that the involution fixes the subgroups $A^{\times}$, $U(\Lambda)$ and $J_{0}(A)$ and so induces an involution on $C(\Lambda)$ given by $\Lambda\alpha]\mapsto[\Lambda\overline{\alpha}]$ (see [CR87, p. 274]).
@@
The involution on $C(\Lambda)$ induced by the involution on $J(A)$ is the standard involution $\ast\colon C(\Lambda)\to C(\Lambda)$, $M]\mapsto-[M^{*}]$.
This gives an alternate way to understand the standard involution on $\widetilde{K}_{0}(\mathbb{Z}G)$. We make further use of this description in the following section.